The Sugar Producers Association wants to estimate the mean yearly sugar consumption. A sample of 25 people reveals the mean yearly consumption to be 27 kg with a sample standard deviation of 9 kg. Assume a normal population. For a 95% confidence interval, what is the critical value needed?
Answer #1

Since the population standard deviation is unknown, and n < 30, we use the students t distribution.

tcritical (2 tailed for a CI) at \alpha = 0.05, degrees of freedom = n - 1 = 25 - 1 = 24 is 2.0639